OpenStudy (anonymous):

Determine if the sequence converges, or diverges. Award Given To Best Answer! An = Sqrt(n)/(1 + sqrt(n))

OpenStudy (henrietepurina):

Write it as a fraction with denominator 1, and then multiply top and bottom by sqrt(n+1) + sqrt(n). This gives: [(n+1) - n] / [ sqrt(n+1) + sqrt(n)] = 1/ [ sqrt(n+1) + sqrt(n)] Now as n gets large the denominator gets large, so the limit as n-> infinity is zero.

OpenStudy (henrietepurina):

Is that ok?

OpenStudy (anonymous):

Yeah, but that doesn't make full sense..Can you explain it more? Or use the equation tab so I can understand it?

OpenStudy (anonymous):

the limit is 1 as n approaches infinity

OpenStudy (anonymous):

I love how I just had the whole thing typed out and then I hit the space bar and it deletes everything. fml.

OpenStudy (anonymous):

Sorry. :P Lol that happens a lot to me too. :P

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty } \frac{ \sqrt{n} }{ 1+\sqrt{n} }\] Here, we will divide by the n term in the DENOMINATOR (only the denominator) with the highest power, which would be the square root of n. \[\lim_{n \rightarrow \infty } \frac{ \frac{ \sqrt{n} }{ \sqrt{n} } }{ \frac{ 1 }{ \sqrt{n} } + \frac{ \sqrt{n} }{ \sqrt{n} } }\] So as n approaches infinity, the 1/sqrt n will get infinitely small, so it's safe to say it would be zero. \[\lim_{n \rightarrow \infty} \frac{ 1 }{ 0+1 } = \frac{ 1 }{ 1 } = 1\] It would converge towards 1.

OpenStudy (anonymous):

@sourwing I can't view the step-by-step solution. :/

OpenStudy (anonymous):

There's your step by step, sir.

OpenStudy (anonymous):

Thank you so much @halorazer :D You're a lifesaver! :)

OpenStudy (anonymous):

Pft, I know. :3

OpenStudy (anonymous):

yes, dividing both top and bottom by sqrt(n) is the key